Bwalya currently has a portfolio of shares giving a return of 21% with a risk of 13%. He is considering a new investment which gives a return of 22% with a risk of 14%. The coefficient of correlation of the new investment with his existing portfolio is +0.2. The new investment will comprise 45% of his enlarged portfolio.

Calculate the portfolio risk and expected return if Bwalya goes ahead with the new investment?


1
Expert's answer
2021-08-25T17:05:44-0400
"r=0.45\times(0.21\times0.13+0.22\times0.14)=0.026"

"d=\frac{45+(22-21)}{21}\times100=219"